how do i graph the lineaur equation like this?
what are the steps?
x+y<10
x-y<-5
and how do i graph any lineaur equation?

Answer by mananth(16949)

x+y <10
you need to graph it.
First find the boundary of the inequality
let
y = -x+10
for x=0, y = 10 (0,10)
For y=0, x=10 (10,10)
Plot the points and draw the line

Now the inequality
x+y<10
y<-x+10
since y is less than something shade the portion below the line

m.ananth@hotmail.ca